OpenStudy (jordanloveangel):

The width of a rectangle is 6 in. less than its length. The perimeter is 68 in. What is the width of the rectangle? i think its 14 inches

TheSmartOne (thesmartone):

Width= Length -6 Perimeter = 2( width + length)

OpenStudy (jordanloveangel):

Length = L Width = L - 6 4L-12 = 68 4L=80 L=20inches W=14 inches

TheSmartOne (thesmartone):

So then perimeter = 2(length- 6+length) 68=2( 2 length -6) 68=4 length -12 4 x length = 80 Length= 20

OpenStudy (jordanloveangel):

so its not 14 its 20 ?

TheSmartOne (thesmartone):

Going back to the first equation width= length -6 width = 20-6 width = 14

TheSmartOne (thesmartone):

It is 14 @jordanloveangel

OpenStudy (jordanloveangel):

oh okay thanx

TheSmartOne (thesmartone):

No problem.